### v1

# AniFaceDiff: High-Fidelity Face Reenactment via Facial Parametric Conditioned Diffusion Models

Abstract
Face reenactment refers to the process of transferring the pose and facial expressions from a reference (driving) video onto a static facial (source) image while maintaining the original identity of the source image. Previous research in this domain has made significant progress by training controllable deep generative models to generate faces based on specific identity, pose and expression conditions. However, the mechanisms used in these methods to control pose and expression often inadvertently introduce identity information from the driving video, while also causing a loss of expression-related details. This paper proposes a new method based on Stable Diffusion [1], called AniFaceDiff, incorporating a new conditioning module for high-fidelity face reenactment. First, we propose an enhanced 2D facial snapshot conditioning approach by facial shape alignment to prevent the inclusion of identity information from the driving video. Then, we introduce an expression adapter conditioning mechanism to address the potential loss of expression-related information. Our approach effectively preserves pose and expression fidelity from the driving video while retaining the identity and fine details of the source image. Through experiments on the VoxCeleb dataset [2], we demonstrate that our method achieves state-of-the-art results in face reenactment, showcasing superior image quality, identity preservation, and expression accuracy, especially for cross-identity scenarios. Considering the ethical concerns surrounding potential misuse, we analyze the implications of our method, evaluate current state-of-the-art deepfake detectors, and identify their shortcomings to guide future research.
1 Introduction
Face reenactment produces a video of a person engaged in conversation by transferring expressions and poses from a reference (driving) video to a source image. This technique finds applications in various fields, including film production, the video game industry, and video conferencing. Two crucial objectives define this task: 1) preserving the identity and background details of the source image; and 2) transferring pose and expression from the driving video onto the static source image. However, achieving a faithful generation that fulfills these criteria remains a significant challenge.
The advancement of deep generative models, including Generative Adversarial Networks (GANs) [3] and diffusion models [4], has facilitated the successful deployment of convenient and high-quality face reenactment techniques. A typical line of the state-of-the-art GANs or diffusion models based methods [5, 6, 7, 8] relies on flow fields to complete the transferring of the pose and the expression. Initially, keypoints of the source image and the driving video are extracted in an unsupervised manner, followed by the estimation of motion flow fields based on these keypoints to capture pose and expression changes. These flow fields are then utilized to warp either the source image or its features. However, the incorporation of estimated flow fields inadvertently introduces identity information from the driving video onto the source image, resulting in a subpar performance in cases of identity mismatch between the source image and the driving video [9]. Moreover, flow fields often cause distortions, particularly in situations involving significant pose and expression variations between the source image and the driving video.
More recent research on face reenactment has highlighted the potential of pretrained 3D parametric face models [10, 11, 12], for representing the semantic meaning of human faces, including expressions, poses, and identities. Methods based on StyleGAN [13], e.g., StyleHeat [14] and HyperReenact [15], utilize parameters or feature maps from 3D face models to adjust the feature map or weights of GAN-generators. However, these approaches are limited by their reliance on finding a latent representation of the source image within the StyleGAN space that can be manipulated. This often leads to inadequate reconstruction of the source image, including both identity information and background details. Diffusion models have also been extended alongside 3D face models [16, 17] owing to their remarkable performance. However, the conditioning mechanisms of these methods still introduce unexpected identity-related information due to the direct use of the driving image to generate the condition. Furthermore, the typically low resolution of 3D face meshes leads to the loss of necessary expression-related information, particularly mid-frequency details [12].
In this work, we propose a method named AniFaceDiff, for face reenactment by incorporating conditional signals to the pretrained text-to-image Stable Diffusion model [1], as shown in Fig. 1. We introduce a new conditioning mechanism to address the limitations of previous methods. Specifically, 1) to avoid introducing unexpected identity-related information from the driving image, we propose a Facial Shape Alignment (FSA) strategy to form spatial-aligned conditions for the diffusion models. This involves extracting pose and expression parameters from the driving video and shape parameters from the source image using the 3D face model (DECA [12]), and then generating 2D face surface normal snapshots with these facial parameters. 2) To mitigate potential expression-related information loss, we introduce an Expression Adapter (EA) that injects the expression embedding into the cross-attention layers of the denoising UNet [18]. This adapter compensates for the information loss in spatial-aligned conditions, thereby enhancing expression fidelity.
In summary, our main contributions are:
1) We propose a face reenactment framework, AniFaceDiff, based on diffusion models with accurate conditioning. Our approach effectively preserves pose and expression fidelity from the driving video simultaneously retaining the identity and background details of the source image in a one-shot setting (solely utilizing a single source image).
2) We design a new conditioning mechanism based on 3D face model, which furnishes both accurate spatial and non-spatial conditions for diffusion models, thereby facilitating faithful reenactment.
3) Experiments on the VoxCeleb [2] benchmark demonstrate that our method achieves state-of-the-art performance in face reenactment tasks. Furthermore, our method has robust generalization capabilities, even when applied to out-of-domain data.
Figure 1: The overview of the proposed method. A VAE [19] encoder and the ReferenceNet extract detailed features from the source image, which are then merged into the Stable Diffusion backbone Denoising UNet via spatial-attention. A CLIP Image Encoder [20] extracts semantic features from the source image, which is then injected into both the ReferenceNet and the Denoising UNet via cross-attention. The pose and expression conditioning is portrayed within the red dashed box. The facial shape of the source image and the pose and expression of the driving video are extracted to form 2D surface normal facial snapshots through Facial Shape Alignment. These snapshots are then encoded and concatenated with the input of the Denoising UNet. The expression of driving frames is further injected into the cross-attention layers of the Denoising UNet by the expression adapter to improve the expression consistency. Temporal-attention aims to improve the temporal consistency across frames. After iterative denoising, the output of the Denoising UNet is decoded into the final animated video by a VAE decoder.
2 Related Work
Neural face reenactment
The development of deep generative models has facilitated convenient and high-quality face reenactment applications. A prevalent approach involves warping the source image or latent representation using deformation fields (e.g., optical flow fields) to convey pose and transfer expressions [21, 22, 5, 6, 9, 7, 23, 24, 25, 26, 27, 8]. For instance, Monkey-Net [22] employs a motion transfer network alongside unsupervised keypoint detection and dense motion prediction to generate animated frames. FOMM [5] enhances Monkey-Net by incorporating motion field computation with a first-order Taylor expansion approximation, encompassing keypoints and affine transformations. Face vid2vid [6] extends FOMM for free-view talking face video generation utilizing a free-view keypoints representation. TPSMM [7] introduces thin-plate spline motion estimation to generate a more flexible optical flow. However, optical flow fields are susceptible to introducing artifacts and blur, especially in instances of large identity or pose mismatches between the source image and driving video. To mitigate this issue, MRAA [9] proposes an Animation Via Disentanglement (AVD) network to separate the control of shape and pose. However, improvements in results are not as significant for objects like faces.
Another line of research exploits the learned prior of pretrained 3D parametric face models. PIRenderer [28] utilizes the parameters of 3D morphable face models (3DMMs) to adjust intermediate results from the warping network. HeadGAN [29] generates faces with warped source features, 3d face shape and audio. StyleHeat [14], based on StyleGAN [13], integrates 3DMMs with StyleGAN by predicting flow fields with 3DMM parameters and warping the feature map from the encoder of GAN inversion. However, these methods based on 3D face models still depend on flow fields and often necessitate additional refinement to attain satisfactory results. In addition, StyleGAN-based methods rely on GAN inversion, which may struggle to achieve high-fidelity reconstruction following extensive edits [15]. HyperReenact [15] employs the feature map of the pretrained DECA encoder to update the weights of the pretrained StyleGAN generator, allowing it to perform well even under large pose variations. PASL [30] proposes pose adapted shape learning for large pose reenactment. Our method aims to enhance the accuracy of both identity preservation and expression.
Image generation with diffusion models
Diffusion models have achieved success across various tasks including unconditional image generation [4, 31, 32], image super-resolution [32], text-to-image generation [33], image-to-image translation [34] and even video generation [35, 36, 37, 38]. Diffusion models possess the capability to capture the full distribution of datasets, making them increasingly popular compared to GANs in recent years. The iterative refinement process also results in the production of diverse and high-quality generated images. Stable Diffusion [1] extends diffusion models into the latent space to significantly reduce computational costs and achieve superior results compared to those in the pixel space.
Conditioning serves as a crucial element in maximizing the potent generative capabilities for various downstream tasks, enabling controllable generation [39, 40, 16, 17, 41, 42, 43]. The two primary conditioning mechanisms for diffusion models involve concatenation and cross-attention [1]. Liu et al. [44] propose an alternative intuitive approach based on [32] and a CLIP-based encoder for semantic diffusion guidance in both text and image conditioning. ControlNet [45] is extensively utilized for spatial-aligned conditioning by updating each layer of the Stable Diffusion backbone UNet via trainable copy and zero convolution layers. IP-adapter [46] integrates image conditions into pretrained text-to-image diffusion models using a decoupled cross-attention mechanism. Our pose and expression conditioning module employs designs similar to those of ControlNet and IP-adapter.
Human image animation with diffusion models
Animating a static image into a temporally consistent video [47, 48, 49, 50, 51] has garnered significant attention among researchers. In human image animation, DisCo [52] introduces human appearance conditioning to Stable Diffusion using a CLIP image encoder and separate ControlNets for background and pose conditions to generate human dancing videos. Animate Anyone [53] employs spatial attention and a ReferenceNet, a copy of UNet, to incorporate detailed information into Stable Diffusion and demonstrate its generalization capability. Champ [54] enhances previous approaches by utilizing a 3D human parametric model with multiple pose conditions. In audio-driven talking head generation, EMO [55] proposes a direct audio-to-video framework leveraging pretrained wave2vec [56]. VLOGGER [57] predicts face and body parameters from audio to render dense masks as conditions for generating talking avatars. We are inspired by previous efforts utilizing the ReferenceNet and spatial-aligned pose conditions due to the similarity of the task. We focus on achieving accurate face reenactment guided by pose and expression while minimizing the impact of identity mismatches.
3 Method
Our method is designed for one-shot face reenactment, which involves generating a realistic video based on the guidance of a single source image and a driving video. The framework of the proposed method is illustrated in Fig. 1. First, we give a brief introduction of the foundational model – Stable Diffusion and the 3D face model (Sec.3.1). Second, we introduce the overall structure of the framework (Sec.3.2). Then, we detail the proposed pose and expression conditioning mechanism (Sec.3.3). Finally, we illustrate the detailed information of training and inference in Sec.3.4
3.1 Preliminaries
Stable Diffusion. Denoising Diffusion Probabilistic Models (DDPM) [4] are a class of generative models that operates by simulating data through the introduction of noise and subsequently denoising it in a progressive manner to generate samples representative of the true data distribution. This step-by-step noising and denoising process endows the model with the ability to generate high-quality images. However, it comes with a high computational resource requirement.
To address this challenge, Stable Diffusion [1] conducts the noising and denoising process in a latent space rather than the pixel space. This enables considerable computational cost savings. The high efficiency allows Stable Diffusion to be pretrained on extremely large-scale datasets (e.g. LAION-5B [58]). Specifically, Stable Diffusion utilizes a pretrained autoencoder to encode the given image $\textbf{x}_{0}$ into a latent representation $\textbf{z}_{0}$ using the encoder. Subsequently, the latent representation can be reconstructed into the pixel space by the decoder. In the training process, Stable Diffusion adds Gaussian noise $\epsilon$ to the latent representation $\textbf{z}_{0}$ at each timestep t by a noise scheduler. The backbone denoising UNet is trained to predict the added noise $\epsilon$. During the inference process, the starting point is usually a randomly selected noise. The trained UNet is utilized to predict the noise under the text condition at each timestep and subsequently conducts the denoising process step by step until a clean image is generated.
3D morphable face models (3DMMs). 3DMMs are parametric models designed to accurately represent facial shape and expression. These models are constructed using dimensionality reduction techniques such as principal component analysis (PCA) and are capable of reconstructing 3D faces based on 2D images. FLAME [11] is a notable example of a 3DMM that relies on standard vertex-based linear blend skinning with blendshapes to generate a mesh with 5023 vertices, which is formulated as:
$M(\mathbf{\beta},\mathbf{\theta},\mathbf{\psi})=W(T_{p}(\mathbf{\beta},\mathbf{\theta},\mathbf{\psi}),\mathbf{J}(\mathbf{\beta}),\mathbf{\theta},\mathcal{W})$ (1)
where $\beta$, $\theta$, and $\psi$ represent parameters of identity, pose, and expression, respectively. $W$ is the blend skinning function that rotates the vertices in $T_{p}$ around joints J smoothed by blendweights $\mathcal{W}$. Detailed information can be found in FLAME [11]. In this paper, we utilize the widely used 3D face model DECA [12] which is capable of estimating parameters, including those of the FLAME model, from a single image. This allows for the reconstruction of a 3D face with detailed facial geometry.
3.2 Framework Architecture
In this section, we illustrate the overall framework of our method, which takes a source image and a driving video as inputs and outputs a reenacted video. We formulate face reenactment as conditional generation, where all the identity, pose, and expression of the generated content are controllable via the corresponding conditioning mechanism. Our method is based on the pretrained Stable Diffusion 1.5 and follows a similar structure to the backbone denoising UNet. First, the source image is encoded into a latent space using a VAE encoder, and its features are extracted by a ReferenceNet. Second, we utilize a CLIP image encoder to extract semantic information, which is crucial for preserving the identity of the source image. This semantic information is injected into the ReferenceNet and Denoising UNet via cross-attention. To effectively introduce pose and expression, we propose a pose and expression conditioning module as elaborated in 3.3. Furthermore, we employ Temporal Modules to generate temporally consistent content as we detail below. The diffusion model performs iterative denoising on the latent space and finally transforms the denoised output back to the pixel space through a VAE decoder to get the generated video.
ReferenceNet. ReferenceNet is a copy of the backbone denoising UNet (without temporal modules) used to extract features at multiple resolutions containing the face and background of the source image. It has been widely utilized [53, 55, 54] to improve the appearance consistency between the reference and the output. These features are then merged into the denoising UNet using a spatial-attention mechanism similar to Animate Anyone [53]. Specifically, the output of each self-attention layer of ReferenceNet ($z_{1}\in\mathbb{R}^{b\times c\times h\times w}$) is repeated $\mathit{f}$ times (the length of the video clip) and concatenated with that of the denoising UNet ($z_{2}\in\mathbb{R}^{b\times c\times f\times h\times w}$). The model then applies self-attention on the concatenated output ($z_{concat}\in\mathbb{R}^{b\times c\times f\times h\times 2w}$) and takes the first half of the feature map as the final output ($z_{out}\in\mathbb{R}^{b\times c\times f\times h\times w}$).
Temporal Module. We employ temporal modules similar to AnimateDiff [59] to improve the temporal consistency across the generated frames. Specifically, the feature map of 3D denoising UNet $z\in\mathbb{R}^{b\times c\times f\times h\times w}$ is firstly reshaped as $z\in\mathbb{R}^{(b\times h\times w)\times f\times c}$. The reshaped feature map is then operated with temporal-attention, which consists of several self-attention blocks along the frame dimension $\mathit{f}$. The temporal module is inserted after the cross-attention layer of each resolution level of denoising UNet via a residual connection.
3.3 Pose and Expression Conditioning
Our pose and expression conditioning module extracts pose and expression from the driving video, and identity-related information from the source image based on a pretrained off-the-shelf 3D face model, DECA. Subsequently, we condition the diffusion model using two components: improved spatial conditioning and non-spatial conditioning (the expression adapter).
Improved Spatial Conditioning with Facial Shape Alignment
We utilize the encoder of DECA to estimate facial parameters from each frame of the driving video directly. Subsequently, we render these parameters into a sequence of 2D facial surface normal snapshots. However, in addition to containing the pose and expression information we require, these snapshots also include identity information reflected by the facial geometry from the driving frames. The identity information essential for our purpose should be extracted from the source image. Any inconsistency in the facial geometry information between the condition and the final output can result in sub-optimal performance, particularly noticeable when there is a significant mismatch in identity between the source image and the driving video. To address this issue, we utilize a facial shape alignment strategy which combines the identity information from the source image with the pose and expression information from the driving frames to generate the refined 2D facial conditions.
Specifically, we obtain shape parameters $\beta$ containing identity information from the source image, and other parameters including pose $\theta$ and expression $\psi$ from the driving frames by the DECA encoder. We then use FLAME to generate the 3D face meshes and render them into surface normal snapshots to eliminate the effects of lighting conditions. The entire process can be represented as:
$n^{1:N}=\mathcal{R}(M(\beta_{s},\theta^{1:N}_{d},\psi^{1:N}_{d}),c^{1:N}_{d}),$ (2)
where $n^{1:N}$ represents the 2D facial conditions, the rendered surface normal snapshots, for frames $1:N$. $\mathcal{R}$ represents the rasterizer (PyTorch3D [60]). M represents FLAME model used to obtain head vertices with facial parameters. $c$ represents the camera information to project the 3D mesh into image space. $\beta_{s}$ is extracted from the source image, while $\theta^{1:N}_{d}$, $\psi^{1:N}_{d}$ and $c^{1:N}_{d}$ are extracted from the driving frames.
After extracting the 2D facial conditions which are spatial-aligned with the output frames, we encode them with the Pose Encoder, a series of convolution layers, similar to Animate Anyone [53], to project the conditions to the same resolution with the noisy input of denoising UNet. Subsequently, we concatenate the spatial-aligned condition with the noisy input and feed it to the denoising UNet.
Expression Adapter
Crucially, the low resolution of the 3D face mesh leads to the loss of mid-frequency details related to expression. Drawing inspiration from IP-Adapter [46] for image prompt conditioning, we introduce an expression adapter conditioning mechanism to address potential information loss resulting from spatial modulation, as illustrated in the lower branch of Fig. 1. Unlike IP-Adapter for conditional image generation, we integrate video-level expression prompts from the driving frames in parallel with the image prompt from the source image. Specifically, we extract the expression embedding $\mathbf{\psi}\in\mathbb{R}^{N\times 50}$ from N driving frames using DECA and project it into the same dimension as the CLIP image embedding. This projected expression embedding $\mathit{p_{\psi}}$ is then injected into the diffusion model through additional cross-attention layers, which can be represented as:
$Z_{new}=Attention(Q,K^{i},V^{i})+\lambda\cdot Attention(Q,K^{\psi},V^{\psi}),$ (3)
where $\mathit{Q}$, $\mathit{K^{i}}$, and $\mathit{V^{i}}$ are the query, key, and value matrices from the original cross-attention layer for CLIP image features, while $\mathit{K^{\psi}}$ and $\mathit{V^{\psi}}$ are from the additional cross-attention layer for expression features. The projected expression embedding $\mathit{p_{\psi}}$ is injected into the cross-attention layer as $\mathit{K^{\psi}=p_{\psi}W_{k}^{\psi}}$ and $\mathit{V^{\psi}=p_{\psi}W_{v}^{\psi}}$, where $\mathit{W_{k}^{\psi}}$ and $\mathit{W_{v}^{\psi}}$ are the corresponding weight matrices. In the 3D denoising UNet for generating a video, the CLIP embedding of the source image is repeated N times to match the length of the expression embedding from the driving frames.
3.4 Training and Inference
The training process comprises two stages. In the first stage, the model is trained to generate a single image under the guidance of a source image and a driving frame. During the training process, the model utilizes same-identity construction as its objective, where both the source image and the driving frame are derived from the same video. DECA is utilized to extract facial parameters and render facial snapshots from the driving frame. The driving frame $\textbf{x}_{0}$ is first encoded into the latent representation $\textbf{z}_{0}$, which is then noised with $\epsilon\sim\mathcal{N}(0,1)$ at timestep t via a defined scheduler as:
$\textbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\textbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ (4)
where $\bar{\alpha}_{t}={\textstyle\prod_{s=0}^{t}}\alpha_{s}$, $\alpha_{t}$ is the variance schedule, and $\textbf{z}_{t}$ is the noisy latent. The model is trained to predict the added noise by denoising function $\epsilon_{\theta}$ with condition features f from the source image and the driving video. The VAE, CLIP image encoder, and DECA are pretrained and kept frozen while the pose encoder, the expression adapter, the ReferenceNet, and the denoising UNet are updated. The training objective of the first stage is similar to that of Stable Diffusion [1]:
$\textbf{L}_{stage1}=\mathbb{E}_{\textit{t},\textbf{f},\epsilon,\textbf{z}_{t}}[\|\epsilon-\epsilon_{\theta}(\textbf{z}_{t},\textit{t},\textbf{f})\|^{2}_{2}]$ (5)
In the second stage, the models are trained to generate a temporally consistent video clip (N frames) with guidance from both a source image and a driving video clip. Pretrained temporal modules [59] are integrated into the denoising UNet. During this stage, only the temporal modules are updated and all other components remain frozen. The training objective of the second stage is formulated as:
$\textbf{L}_{stage2}=\mathbb{E}_{\textit{t},\textbf{f}^{1:N},\epsilon,\textbf{z}_{t}^{1:N}}[\|\epsilon-\epsilon_{\theta}(\textbf{z}_{t}^{1:N},\textit{t},\textbf{f}^{1:N})\|^{2}_{2}]$ (6)
At inference, we utilize DECA to extract information from both the source image and the driving video. Leveraging a facial shape alignment strategy, we refine the facial snapshot sequence. Begining with noises $\textbf{z}_{t}^{1:N}$ sampled from a Gaussian distribution in the latent space, the model conduct denoising process to $\textbf{z}_{t-1}^{1:N}$ by predicting noise using the denoising UNet with the condition $\textbf{f}^{1:N}$ and iteratively to $\textbf{z}_{0}^{1:N}$ via a defined sampling process. Finally, the denoised outputs $\textbf{z}_{0}^{1:N}$ are projected back to the pixel level $\textbf{x}_{0}^{1:N}$ to obtain the generated video via the VAE decoder.
4 Experiments
In our experiments, we utilized the VoxCeleb dataset [2] after preprocessing as described in [5]. This dataset comprises 19,522 and 525 utterance sequences, spanning 453 and 13 identities, with 3,354 and 90 videos in the training and test sets, respectively.
Training was conducted on $256\times 256$ images using 2 NVIDIA A100 GPUs. In the first stage, we trained for over 100,000 steps with a batch size of 32. Subsequently, in the second stage, we trained for over 20,000 steps using a batch size of 8 for a video clip consisting of 24 frames. All models were optimized using the Adam optimizer with a learning rate of $1\times 10^{-5}$.
4.1 Comparisons
We compare our method with four state-of-the-art methods, namely FOMM [5], Face vid2vid [6], TPSMM [7], and HyperReenact [15]. Results for the baseline methods are generated using their provided code and pretrained weights, with the exception of Face vid2vid, for which we utilize unofficial code [^1] https://github.com/zhanglonghao1992/One-Shot_Free-View_Neural_Talking_Head_Synthesis due to the absence of the official version. We assess our method and the baseline approaches using four distinct metrics. First, we utilize the Fréchet-Inception Distance (FID) to evaluate image quality by measuring the distribution discrepancy between the generated data and the test set of VoxCeleb. Second, we employ the identity preservation cosine similarity (CSIM), following the approach outlined in Encoding in Style [61], to quantify the preservation of identity. Third, we measure the accuracy of pose and expression transfer using the Average Pose Distance (APD) and the Average Expression Distance (AED). CSIM is computed based on the source image and the outputs, while APD and AED are calculated based on the driving video and the outputs.
Quantitative Comparisons
Table 1: Quantitative results on VoxCeleb. The best results are shown in bold, and the second-best results are shown in underline
. Same-identity Reconstruction Cross-identity Reenactment FID CSIM APD AED FID CSIM APD AED User Pref. (%) FOMM [5] 21.91 0.83 0.011 0.092 41.93 0.51 0.028 0.210 10.25 Face vide2vid [6] 18.57 0.84 0.012 0.092 35.21 0.56 0.034 0.238 9.50 TPSMM [7] 20.18 0.84 0.009 0.081 38.91 0.52 0.024 0.191 16.50 HyperReenact [15] 72.72 0.48 0.011 0.098 73.35 0.40 0.017 0.181 20.75 Ours 17.21 0.75 0.011 0.091 32.58 0.57 0.022 0.174 43.00
We evaluate our method and the baselines on the test set of VoxCeleb. In the evaluation, higher scores indicate better performance for CSIM, while lower scores are preferable for all other metrics. We conduct one-shot face reenactment on two tasks: same-identity reconstruction, which mirrors the training process, and the more challenging primary task of cross-identity reenactment.
For the same-identity reconstruction task, we randomly selected 100 videos from the VoxCeleb test set, using the first frame as the source image and considering the remaining frames as the driving frames. As shown in Table 1, our method achieves the best image quality (FID) and ranks second for pose (APD) and expression accuracy (AED). TPSMM outperforms all other methods in terms of CSIM, APD, and AED. All warping-based methods (FOMM, Face vid2vid, and TPSMM) perform well on this proxy task when there is no identity variance and usually a small pose difference between the source image and the driving video.
Even though current methods perform well on same-identity reconstruction, the main task (cross-identity reenactment), where the source image and driving video are from different identities, introduces substantial identity mismatches and pose variations, demanding higher generalization capabilities from models. Similar to previous work, we randomly select 100 image-video pairs from the VoxCeleb test set. Table.1 illustrates a significant performance drop for each method when transitioning from same-identity reconstruction to cross-identity reenactment. As observed in previous SOTA methods, there is a trade-off between identity consistency (CSIM) and expression accuracy (AED). The more modifications related to expression made to the original source identity, the less the generated image will resemble the original, as assessed by face recognition models. Despite this trade-off, Our method (AniFaceDiff) achieves SOTA results on both CSIM and AED. Additionally, we outperform all four baseline methods in terms of image quality (FID) and secure the second position in terms of pose accuracy (APD).
Qualitative Comparisons
Figure 2: Qualitative comparison with SOTA methods on cross-identity reenactment. Our method achieves high identity preservation and expression accuracy. (All photorealistic human images presented in this paper are virtual individuals generated by [33] and do not exist in the real world.)
Figure 3: Qualitative comparison with SOTA methods on cross-identity reenactment in the presence of significant pose variations between the source and driving images. Our method accurately captures the source identity and faithfully transfers the target pose and expression, even under significant pose variations.
We present a few representative examples of cross-identity reenactment in Fig. 2. Additional examples can be found in Appendix A.2. Overall, our method generates high quality images compared to all other state-of-the-art methods while preserving both source identity and pose and expression of the driving frame. Better preservation of source identity is especially evident in the first and second rows, where other methods show noticeable influence from the driving identity. In contrast, our method maintains the source identity with greater accuracy. Furthermore, our approach excels at maintaining the expressions and poses of the driving videos, as seen in the third and fourth rows, where our method consistently captures expression-related details such as wrinkles, brow furrows and lip curvature more accurately than other state-of-the-art methods.
Fig.3 demonstrates a more challenging scenario where there are significant pose differences between the source and driving images. FOMM, Face vid2vid, and TPSMM tend to produce blurry outputs with severe artifacts and distortions due to their reliance on the optical flow field. Additionally, their inadequate inpainting capabilities result in blurry and unrealistic regions when generating areas not present in the source image. HyperReenact, which does not rely on the optical flow field, generates more fine-grained results. However, it still suffers from a loss of identity and detailed information, such as facial wrinkles, hair texture, and lighting. This is likely due to the limitations of GAN inversion and the use of a pretrained ArcFace [62] for extracting identity information from the source image. In contrast, our method excels in preserving detailed information from the source image, while accurately maintaining pose and expression from the driving video due to the designed conditioning module. Overall, our method produces the cleanest and most high-fidelity results among all the state-of-the-art methods, even when the source face is partially occluded (row 3) or when there is a large distance difference from the camera between the source and driving images (row 4).
We note that the current metrics might not fully capture the qualitative differences observed in the results. Hence, we conducted a user study involving 20 participants to further assess performance based on human visual experience. We presented 20 randomly selected source-driving image pairs: 5 for same-identity reconstruction and 15 for cross-identity reenactment, as cross-identity reenactment is the primary task. We provided users with three evaluation criteria similar to those used by HyperReenact [15]: 1) image quality, 2) identity and nuanced detail preservation from the source image, and 3) pose and expression preservation from the driving image. As shown in Table 1, our method significantly outperforms all other SOTA methods based on user preference.
Out-of-domain Data
Figure 4: Qualitative comparison with SOTA methods on out-of-domain (non-photorealistic) face reenactment. Our method can generate high-quality images even without being trained on such types of data. (All out-of-domain facial images are generated by [63])
We also compare our method with these SOTA methods outside the training distribution to evaluate generalization capabilities. As shown in Fig. 4, The first and second rows demonstrate out-of-domain (non-photorealistic) source images reenacted by out-of-domain driving images. The third and fourth rows demonstrate out-of-domain source images reenacted by in-domain (photorealistic) driving images. Our method exhibits strong generalization capabilities on out-of-domain data. More results can be found in Appendix A.3.
4.2 Ablation Study
Table 2: Quantitative comparison with or without Facial Shape Alignment and the Expression Adapter. The addition of both components allows for a better trade-off between identity consistency vs expression fidelity. We find that combining both provides the best qualitative and quantitative results.
. Same-identity Reconstruction Cross-identity Reenactment FSA EA CSIM APD AED CSIM APD AED - - 0.75 0.011 0.098 0.52 0.021 0.182 - - - - 0.59 0.022 0.206 - 0.75 0.011 0.091 0.51 0.021 0.158 - - - 0.57 0.022 0.174
Figure 5: Qualitative comparison on cross-identity reenactment with or without Facial Shape Alignment and the Expression Adapter. FSA improves facial generation quality, while EA maintains expression consistency. Ensembling both provides the best overall consistency. These observations are consistent with the quantitative results in Table 2.
We conduct an ablation study on the proposed 1) Facial Shape Alignment (FSA) and 2) Expression Adapter (EA). Facial Shape Alignment is used exclusively for cross-identity reenactment. Our baseline method directly conditions the model with 2D facial snapshots extracted from the driving video, without the facial shape alignment and the expression adapter. As illustrated in Table 2, adding facial shape alignment significantly improves identity preservation (CSIM). Facial shape alignment uses the shape embedding from the source image to render the 2D facial snapshots, thereby enhancing identity preservation. Without facial shape alignment, the output maintains the shape from the driving image rather than the source image, as shown in Fig. 5. The expression adapter significantly enhances expression accuracy (AED) in both same-identity reconstruction and cross-identity reenactment. Methods incorporating the expression adapter can provide more expression-related details, such as wrinkles, and achieve greater accuracy in certain areas, such as the mouth, as depicted in Fig. 5. Combining both FSA and EA can effectively handle the trade-off between identity consistency and expression accuracy.
5 Ethical Considerations
We are committed to positive virtual face reenactment applications. All photorealistic human portraits presented in this paper are virtual and do not exist in reality. Acknowledging the broad applicability of face reenactment, we are also aware of its potential for malicious misuse, such as in deepfake fraud. To address this concern, we plan to implement stringent access control mechanisms to limit access permissions exclusively for research purposes. Additionally, we evaluate existing SOTA deepfake detectors (see Appendix A.1) and conduct an analysis to identify potential shortcomings. This will serve as a foundation for future research aimed at mitigating weaknesses of existing detectors and providing benchmarking data for future approaches.
Appendix A Appendix / supplemental material
A.1 On deepfake detection and potential for misuse
While any face reenactment technique can also be misused for deepfake generation, it is clearly an important area of active research with many ethical applications in movies, journalism, animation and virtual reality, to name a few. This is especially apparent in emergent capabilities in techniques such as ours, as improved results in the primary domain (real human face animation) can lead to significant improvements in cross-domain or out of domain animation transfer (human to 3D, human to cartoon, cartoon to 3D, cartoon to cartoon etc.) However, to mitigate the deepfake generation potential of our approach, we perform a detailed and comprehensive analysis of state of the art detection techniques on our method, and report their zero shot performance on our data. First, this contributes to the understanding of the generalization capabilities of existing deepfake detectors. Secondly, it highlights potential for improving the training datasets of such detectors, especially with videos based on paradigms using diffusion models as the generative core. This work opens up the possibility of creating a large scale dataset of high fidelity diffusion based face reenactment videos, addressing the issue of lack of such data in the deepfake detection domain.
Deepfake Autoencoder (DFAE, e.g. DeepFaceLab [^2] https://github.com/iperov/DeepFaceLab) and other face swapping methods can produce highly realistic videos that are challenging to detect using existing detection algorithms. This is evident in the results of detection algorithms on the DFDC [64] dataset. Additionally, videos generated by face reenactment methods such as Neural Texture [65] and Face2Face [66] are also difficult to detect using off-the-shelf detectors (details in [67]). It is worth noting that these methods do not offer any alternative means of detecting videos generated by them, which encourages generation of defamatory and harmful deepfakes.
Model Frame-AUC Video-AUC
Xception [68] 85.03 87.73
Capsule [69] 79.43 79.06
EfficientNet-B4 [70] 81.20 85.96
FFD [71] 84.79 82.59
SRM [72] 85.37 88.72
RECCE [73] 87.42 85.00
CORE [74] 89.32 87.12
CADDM [75] 86.82 95.99
Average 84.92 86.52
Table 3: Detection frame-level and video-level AUC from the state-of-the-art deepfake detectors with respective pretrained weights on videos generated by our reenactment method AniFaceDiff.
We emphasize the significance of ensuring that these methods are used ethically and that they can be detected with existing deepfake detectors; this would invalidate videos generated with malicious intent. We evaluated videos generated with our method on pretrained, state-of-the-art deepfake detection methods. Table 3 presents both frame and video-level detection results. Although video-level detection is more common, we include frame-level detection due to the possibility of temporally partial deepfakes [67] i.e. having both real and fake segments in the same video. Our evaluation encompasses multiple types of detectors including naive detectors [68, 70], a frequency-level detector [72] and spatial detectors [69, 71, 73, 74, 75]. We notice that videos generated with AniFaceDiff had high detection rate by the tested detectors. Frame-level detection is more challenging since the detection performance depends on the prediction of each frame, rather than a majority voting across the frames in a video. While the results are promising, there is room for improvement. Further investigation into diffusion-based reenactment methods could further enhance detection rates.
This assessment aims to pave the way for future studies on detection. Additionally, we provide access to researchers from deepfake detection in order to facilitate the generation of higher quality deepfake datasets, particularly those generated using state of the art face reenactment diffusion models (such as our work) - a current gap in the deepfake literature.
A.2 More results on cross-identity reenactment
Figure 6: Additional qualitative comparison with SOTA methods on cross-identity reenactment.
A.3 More results on out-of-domain reenactment
Figure 7: Additional qualitative comparison with SOTA methods on out-of-domain reenactment.
A.4 Limitations
Although our method provides high-fidelity results with excellent image quality, there are still some limitations. First, noticeable flickering occurs when generating videos. Second, the inadvertent generation of hands affects the quality of the generated content. Future studies can focus on addressing these issues.

### v2

@@
-AniFaceDiff: High-Fidelity Face Reenactment via Facial Parametric Conditioned Diffusion Models
+AniFaceDiff: Animating Stylized Avatars via Parametric Conditioned Diffusion Models
 Abstract
-Face reenactment refers to the process of transferring the pose and facial expressions from a reference (driving) video onto a static facial (source) image while maintaining the original identity of the source image.
-Previous research in this domain has made significant progress by training controllable deep generative models to generate faces based on specific identity, pose and expression conditions.
-However, the mechanisms used in these methods to control pose and expression often inadvertently introduce identity information from the driving video, while also causing a loss of expression-related details.
-This paper proposes a new method based on Stable Diffusion [1], called AniFaceDiff, incorporating a new conditioning module for high-fidelity face reenactment.
-First, we propose an enhanced 2D facial snapshot conditioning approach by facial shape alignment to prevent the inclusion of identity information from the driving video.
-Then, we introduce an expression adapter conditioning mechanism to address the potential loss of expression-related information.
-Our approach effectively preserves pose and expression fidelity from the driving video while retaining the identity and fine details of the source image.
-Through experiments on the VoxCeleb dataset [2], we demonstrate that our method achieves state-of-the-art results in face reenactment, showcasing superior image quality, identity preservation, and expression accuracy, especially for cross-identity scenarios.
-Considering the ethical concerns surrounding potential misuse, we analyze the implications of our method, evaluate current state-of-the-art deepfake detectors, and identify their shortcomings to guide future research.
+Animating stylized avatars with dynamic poses and expressions has attracted increasing attention for its broad range of applications.
+Previous research has made significant progress by training controllable generative models to synthesize animations based on reference characteristics, pose, and expression conditions.
+However, the mechanisms used in these methods to control pose and expression often inadvertently introduce unintended features from the target motion, while also causing a loss of expression-related details, particularly when applied to stylized animation.
+This paper proposes a new method based on Stable Diffusion [1], called AniFaceDiff, incorporating a new conditioning module for animating stylized avatars.
+First, we propose a refined spatial conditioning approach by Facial Alignment to prevent the inclusion of identity characteristics from the target motion.
+Then, we introduce an Expression Adapter that incorporates additional cross-attention layers to address the potential loss of expression-related information.
+Our approach effectively preserves pose and expression from the target video while maintaining input image consistency.
+Extensive experiments demonstrate that our method achieves state-of-the-art results, showcasing superior image quality, preservation of reference features, and expression accuracy, particularly for out-of-domain animation across diverse styles, highlighting its versatility and strong generalization capabilities.
+This work aims to enhance the quality of virtual stylized animation for positive applications.
+To promote responsible use in virtual environments, we contribute to the advancement of detection for generative content by evaluating state-of-the-art detectors, highlighting potential areas for improvement, and suggesting solutions.
+Figure 1: Stylized animation results produced by our method.
+In the visualizations, the first columns are reference images and the following three columns are animated results.
+Our method generates high-quality results using various reference and target motion images that differ in pose, expression, and even domain.
+In particular, it effectively preserves the reference appearance and style with the pose and expression of the target motion, even under significant pose variations.
+Furthermore, it demonstrates a strong capability to handle avatars across a wide range of styles and domains.
+We aim on virtual stylized character animation.
+All characters in this paper are generated by [2] and do not correspond to real individuals.
+More examples can be found in Fig.
+6.
 1 Introduction
-Face reenactment produces a video of a person engaged in conversation by transferring expressions and poses from a reference (driving) video to a source image.
-This technique finds applications in various fields, including film production, the video game industry, and video conferencing.
-Two crucial objectives define this task: 1) preserving the identity and background details of the source image; and 2) transferring pose and expression from the driving video onto the static source image.
-However, achieving a faithful generation that fulfills these criteria remains a significant challenge.
-The advancement of deep generative models, including Generative Adversarial Networks (GANs) [3] and diffusion models [4], has facilitated the successful deployment of convenient and high-quality face reenactment techniques.
+The animation of stylized avatars refers to the process of animating static, stylized characters by poses and expressions derived from target motion.
+This technique has a wide range of applications for film and animation production, video conferencing, customer service, and privacy protection.
+For instance, it can be utilized to safeguard patients’ appearance privacy in telemedicine and to enhance engagement in virtual learning through artistic representations.
+Two crucial objectives define this task: 1) preserving the appearance and background details of the stylized reference image; and 2) accurately reflecting the pose and expressions from the target motion.
+However, achieving high-quality stylized animation that meets these objectives remains a significant challenge.
+The advancement of deep generative models, including Generative Adversarial Networks (GANs) [3] and diffusion models [4], has facilitated the success of high-quality avatar animation.
 A typical line of the state-of-the-art GANs or diffusion models based methods [5, 6, 7, 8] relies on flow fields to complete the transferring of the pose and the expression.
-Initially, keypoints of the source image and the driving video are extracted in an unsupervised manner, followed by the estimation of motion flow fields based on these keypoints to capture pose and expression changes.
-These flow fields are then utilized to warp either the source image or its features.
-However, the incorporation of estimated flow fields inadvertently introduces identity information from the driving video onto the source image, resulting in a subpar performance in cases of identity mismatch between the source image and the driving video [9].
-Moreover, flow fields often cause distortions, particularly in situations involving significant pose and expression variations between the source image and the driving video.
-More recent research on face reenactment has highlighted the potential of pretrained 3D parametric face models [10, 11, 12], for representing the semantic meaning of human faces, including expressions, poses, and identities.
-Methods based on StyleGAN [13], e.g., StyleHeat [14] and HyperReenact [15], utilize parameters or feature maps from 3D face models to adjust the feature map or weights of GAN-generators.
-However, these approaches are limited by their reliance on finding a latent representation of the source image within the StyleGAN space that can be manipulated.
-This often leads to inadequate reconstruction of the source image, including both identity information and background details.
-Diffusion models have also been extended alongside 3D face models [16, 17] owing to their remarkable performance.
-However, the conditioning mechanisms of these methods still introduce unexpected identity-related information due to the direct use of the driving image to generate the condition.
-Furthermore, the typically low resolution of 3D face meshes leads to the loss of necessary expression-related information, particularly mid-frequency details [12].
-In this work, we propose a method named AniFaceDiff, for face reenactment by incorporating conditional signals to the pretrained text-to-image Stable Diffusion model [1], as shown in Fig.
-1.
+Initially, keypoints of the inputs are extracted in an unsupervised manner, followed by the estimation of motion flow fields based on these keypoints to capture pose and expression changes.
+These flow fields are then utilized to warp either the reference image or its features.
+However, the incorporation of estimated flow fields inadvertently introduces identity characteristics from the motion video onto the reference image, resulting in a subpar performance in cases of identity mismatch between the reference image and the target video [9].
+This issue is particularly pronounced in stylized animation, where the appearance inconsistency can be more significant, further exacerbating the problem.
+Moreover, flow fields often cause distortions, particularly in situations involving significant pose and expression variations between the reference image and the target motion.
+More recent research has highlighted the potential of pretrained 3D Morphable Face Models (3DMMs) [10, 11, 12], for representing the semantic meaning of faces, including expressions, poses, and other features.
+Methods based on StyleGAN [13], such as StyleHeat [14], utilize parameters or feature maps from 3D morphable models to adjust the feature map or weights of GAN-generators.
+However, these approaches are limited by their reliance on finding a latent representation of the reference image within the StyleGAN space that can be manipulated.
+This often leads to inadequate reconstruction of the reference image, including both appearance and background details.
+Diffusion models have also been extended alongside 3D morphable models [15, 16] owing to their remarkable performance.
+However, the conditioning mechanisms of these methods still introduce unexpected identity-related information due to the direct use of the target motion to generate the condition.
+Furthermore, the typically low resolution of 3D face meshes leads to the loss of necessary expression-related information, particularly mid-frequency details [12], which are critical for achieving expressive animation in stylized avatars.
+In this work, we propose a method named AniFaceDiff, for stylized and diverse avatar animation by incorporating conditional signals to the pretrained text-to-image Stable Diffusion model [1].
 We introduce a new conditioning mechanism to address the limitations of previous methods.
-Specifically, 1) to avoid introducing unexpected identity-related information from the driving image, we propose a Facial Shape Alignment (FSA) strategy to form spatial-aligned conditions for the diffusion models.
-This involves extracting pose and expression parameters from the driving video and shape parameters from the source image using the 3D face model (DECA [12]), and then generating 2D face surface normal snapshots with these facial parameters.
-2) To mitigate potential expression-related information loss, we introduce an Expression Adapter (EA) that injects the expression embedding into the cross-attention layers of the denoising UNet [18].
-This adapter compensates for the information loss in spatial-aligned conditions, thereby enhancing expression fidelity.
+Specifically, 1) to avoid introducing unexpected identity-related information from the target motion, we propose a Facial Alignment (FA) strategy to form spatial-aligned conditions for the diffusion models.
+This process involves extracting pose and expression parameters from the target video, as well as shape parameters from the reference image, using the 3D morphable model (DECA [12]).
+These parameters are used to generate 2D surface normal snapshots, which are then encoded as part of the input for the denoising UNet.
+[17].
+2) To compensate for the expression-related information loss in spatial-aligned conditions, we introduce an Expression Adapter (EA) that integrates the expression embedding from the target video into the denoising UNet using additional cross-attention layers, which are combined with the cross-attention layers for the CLIP embedding [18] of the reference image.
+As shown in Fig.
+1, our results demonstrate the effectiveness of our methods in handling avatars with diverse styles and generating high-fidelity outputs.
+Furthermore, we aim to improve virtual stylized animation for positive applications.
+To promote responsible development, we contribute to the improvement of detection for generative content, reinforcing safeguards against potential misuse.
 In summary, our main contributions are:
-1) We propose a face reenactment framework, AniFaceDiff, based on diffusion models with accurate conditioning.
-Our approach effectively preserves pose and expression fidelity from the driving video simultaneously retaining the identity and background details of the source image in a one-shot setting (solely utilizing a single source image).
-2) We design a new conditioning mechanism based on 3D face model, which furnishes both accurate spatial and non-spatial conditions for diffusion models, thereby facilitating faithful reenactment.
-3) Experiments on the VoxCeleb [2] benchmark demonstrate that our method achieves state-of-the-art performance in face reenactment tasks.
-Furthermore, our method has robust generalization capabilities, even when applied to out-of-domain data.
+1) We propose a framework based on diffusion models, incorporating an accurate conditioning approach for animation across diverse styles.
+Our approach effectively preserves pose and expression fidelity from the target motion while maintaining the appearance and background details of the stylized reference image.
+2) We design a new conditioning mechanism based on 3D morphable models, which furnishes both accurate spatial and non-spatial conditions for diffusion models, thereby facilitating faithful animation.
+3) Extensive experiments demonstrate that our method achieves state-of-the-art performance in both out-of-domain and within-domain animation.
+4) We evaluate the generalization capabilities of state-of-the-art synthetic face detectors, helping to identify potential weaknesses and propose possible solutions.
 Figure 2: The overview of the proposed method.
-A VAE [19] encoder and the ReferenceNet extract detailed features from the source image, which are then merged into the Stable Diffusion backbone Denoising UNet via spatial-attention.
-A CLIP Image Encoder [20] extracts semantic features from the source image, which is then injected into both the ReferenceNet and the Denoising UNet via cross-attention.
-The pose and expression conditioning is portrayed within the red dashed box.
-The facial shape of the source image and the pose and expression of the driving video are extracted to form 2D surface normal facial snapshots through Facial Shape Alignment.
-These snapshots are then encoded and concatenated with the input of the Denoising UNet.
-The expression of driving frames is further injected into the cross-attention layers of the Denoising UNet by the expression adapter to improve the expression consistency.
-Temporal-attention aims to improve the temporal consistency across frames.
-After iterative denoising, the output of the Denoising UNet is decoded into the final animated video by a VAE decoder.
+A reference image and a target video serve as inputs.
+The expression from the target frames and the CLIP reference embedding are integrated into the decoupled cross-attention layers of the Denoising UNet via the expression adapter, enhancing expression consistency while preserving the semantic information from the reference image.
+The head shape from the reference image, along with the pose and expression from the target motion, are extracted and scaled to generate 2D surface normal snapshots through Facial Alignment.
+These snapshots are subsequently encoded and incorporated into the input of the Denoising UNet.
+The ReferenceNet plays a key role in preserving detailed features from the reference image, while temporal attention focuses on maintaining temporal consistency across frames.
+After iterative denoising, the final output from the Denoising UNet is decoded into the resulting animated video.
 2 Related Work
-Neural face reenactment
-The development of deep generative models has facilitated convenient and high-quality face reenactment applications.
-A prevalent approach involves warping the source image or latent representation using deformation fields (e.g., optical flow fields) to convey pose and transfer expressions [21, 22, 5, 6, 9, 7, 23, 24, 25, 26, 27, 8].
-For instance, Monkey-Net [22] employs a motion transfer network alongside unsupervised keypoint detection and dense motion prediction to generate animated frames.
-FOMM [5] enhances Monkey-Net by incorporating motion field computation with a first-order Taylor expansion approximation, encompassing keypoints and affine transformations.
-Face vid2vid [6] extends FOMM for free-view talking face video generation utilizing a free-view keypoints representation.
+Motion-guided avatar animation
+The development of deep generative models has facilitated high-quality creation of avatars.
+A prevalent approach involves warping the reference image or latent representation using deformation fields (e.g., optical flow fields) to convey pose and transfer expressions [19, 5, 6, 9, 7, 20, 21, 22, 23, 8].
+For instance, FOMM [5] enhances Monkey-Net by incorporating motion field computation with a first-order Taylor expansion approximation, encompassing keypoints and affine transformations.
+Face vid2vid [6] extends FOMM for free-view talking head video generation utilizing a free-view keypoints representation.
 TPSMM [7] introduces thin-plate spline motion estimation to generate a more flexible optical flow.
-However, optical flow fields are susceptible to introducing artifacts and blur, especially in instances of large identity or pose mismatches between the source image and driving video.
+However, optical flow fields are susceptible to introducing artifacts and blur, especially in instances of large pose mismatches between the reference image and target video.
 To mitigate this issue, MRAA [9] proposes an Animation Via Disentanglement (AVD) network to separate the control of shape and pose.
 However, improvements in results are not as significant for objects like faces.
-Another line of research exploits the learned prior of pretrained 3D parametric face models.
-PIRenderer [28] utilizes the parameters of 3D morphable face models (3DMMs) to adjust intermediate results from the warping network.
-HeadGAN [29] generates faces with warped source features, 3d face shape and audio.
+Another line of research exploits the learned prior of pretrained 3D parametric morphable models.
+PIRenderer [24] utilizes the parameters of 3D morphable models to adjust intermediate results from the warping network.
+HeadGAN [25] generates heads with warped reference features, 3d face shape and audio.
 StyleHeat [14], based on StyleGAN [13], integrates 3DMMs with StyleGAN by predicting flow fields with 3DMM parameters and warping the feature map from the encoder of GAN inversion.
-However, these methods based on 3D face models still depend on flow fields and often necessitate additional refinement to attain satisfactory results.
+However, these methods based on 3D morphable models still depend on flow fields and often necessitate additional refinement to attain satisfactory results.
 In addition, StyleGAN-based methods rely on GAN inversion, which may struggle to achieve high-fidelity reconstruction following extensive edits [26].
 HyperReenact [26] employs the feature map of the pretrained DECA encoder to update the weights of the pretrained StyleGAN generator, allowing it to perform well even under large pose variations.
-PASL [30] proposes pose adapted shape learning for large pose reenactment.
-Our method aims to enhance the accuracy of both identity preservation and expression.
+PASL [27], on the other hand, employs pose-adapted shape learning specifically designed for handling significant pose variations.
+Our method builds on diffusion models and aims to enhance both the preservation of stylized appearance and the accuracy of expression through a carefully designed conditioning module, particularly for avatars with diverse styles.
 Image generation with diffusion models
@@
 IP-Adapter [43] integrates image conditions into pretrained text-to-image diffusion models using a decoupled cross-attention mechanism.
-Our pose and expression conditioning module employs designs similar to those of ControlNet and IP-adapter.
-Human image animation with diffusion models
+Our pose and expression conditioning module is inspired by ControlNet and IP-Adapter.
+Character animation with diffusion models
 Animating a static image into a temporally consistent video [44, 45, 46, 47, 48] has garnered significant attention among researchers.
-In human image animation, DisCo [52] introduces human appearance conditioning to Stable Diffusion using a CLIP image encoder and separate ControlNets for background and pose conditions to generate human dancing videos.
+DisCo [49] introduces appearance conditioning to Stable Diffusion using a CLIP image encoder and separate ControlNets for background and pose conditions to generate dancing videos.
 Animate Anyone [50] employs spatial attention and a ReferenceNet, a copy of UNet, to incorporate detailed information into Stable Diffusion and demonstrate its generalization capability.
@@
 We are inspired by previous efforts utilizing the ReferenceNet and spatial-aligned pose conditions due to the similarity of the task.
-We focus on achieving accurate face reenactment guided by pose and expression while minimizing the impact of identity mismatches.
+Our approach focuses on stylized avatar animation using diffusion models, minimizing identity mismatches through Facial Alignment with shape and scale alignment, and mitigating expression loss using the Expression Adapter.
 3 Method
-Our method is designed for one-shot face reenactment, which involves generating a realistic video based on the guidance of a single source image and a driving video.
+Our method is designed for one-shot animation across diverse styles, which involves generating a video based on the guidance of a single reference image and target motion.
 The framework of the proposed method is illustrated in Fig.
-1.
-First, we give a brief introduction of the foundational model – Stable Diffusion and the 3D face model (Sec.3.1).
+2.
+First, we give a brief introduction of the foundational model – Stable Diffusion and the 3D morphable models (Sec.3.1).
 Second, we introduce the overall structure of the framework (Sec.3.2).
@@
 3D morphable face models (3DMMs).
-3DMMs are parametric models designed to accurately represent facial shape and expression.
+3DMMs are parametric models designed to accurately represent head shape and expression.
 These models are constructed using dimensionality reduction techniques such as principal component analysis (PCA) and are capable of reconstructing 3D faces based on 2D images.
@@
 $M(\mathbf{\beta},\mathbf{\theta},\mathbf{\psi})=W(T_{p}(\mathbf{\beta},\mathbf{\theta},\mathbf{\psi}),\mathbf{J}(\mathbf{\beta}),\mathbf{\theta},\mathcal{W})$ (1)
-where $\beta$, $\theta$, and $\psi$ represent parameters of identity, pose, and expression, respectively.
+where $\beta$, $\theta$, and $\psi$ represent parameters of shape, pose, and expression, respectively.
 $W$ is the blend skinning function that rotates the vertices in $T_{p}$ around joints J smoothed by blendweights $\mathcal{W}$.
 Detailed information can be found in FLAME [11].
-In this paper, we utilize the widely used 3D face model DECA [12] which is capable of estimating parameters, including those of the FLAME model, from a single image.
-This allows for the reconstruction of a 3D face with detailed facial geometry.
+In this paper, we utilize the widely used 3D morphable model DECA [12] which is capable of estimating parameters, including those of the FLAME model, from a single image.
+This allows for the reconstruction of a 3D head with detailed head geometry.
 3.2 Framework Architecture
-In this section, we illustrate the overall framework of our method, which takes a source image and a driving video as inputs and outputs a reenacted video.
-We formulate face reenactment as conditional generation, where all the identity, pose, and expression of the generated content are controllable via the corresponding conditioning mechanism.
+In this section, we illustrate the overall framework of our method, which takes a reference image and a target video as inputs and outputs an animated video.
+We formulate avatar animation as conditional generation, where all the appearance, pose, and expression of the generated content are controllable via the corresponding conditioning mechanism.
 Our method is based on the pretrained Stable Diffusion 1.5 and follows a similar structure to the backbone denoising UNet.
-First, the source image is encoded into a latent space using a VAE encoder, and its features are extracted by a ReferenceNet.
-Second, we utilize a CLIP image encoder to extract semantic information, which is crucial for preserving the identity of the source image.
+First, the reference image is encoded into a latent space using a VAE encoder [56], and its features are extracted by a ReferenceNet.
+Second, we utilize a CLIP image encoder to extract semantic information, which is crucial for preserving the characteristics of the reference image.
 This semantic information is injected into the ReferenceNet and Denoising UNet via cross-attention.
@@
 ReferenceNet.
-ReferenceNet is a copy of the backbone denoising UNet (without temporal modules) used to extract features at multiple resolutions containing the face and background of the source image.
+ReferenceNet is a copy of the backbone denoising UNet (without temporal modules) used to extract features at multiple resolutions containing the head, style, and background of the reference image.
 It has been widely utilized [50, 52, 51] to improve the appearance consistency between the reference and the output.
@@
 3.3 Pose and Expression Conditioning
-Our pose and expression conditioning module extracts pose and expression from the driving video, and identity-related information from the source image based on a pretrained off-the-shelf 3D face model, DECA.
+Our pose and expression conditioning module extracts pose and expression from the target video, and identity-related information from the reference image based on a pretrained off-the-shelf 3D morphable model, DECA.
 Subsequently, we condition the diffusion model using two components: improved spatial conditioning and non-spatial conditioning (the expression adapter).
-Improved Spatial Conditioning with Facial Shape Alignment
-We utilize the encoder of DECA to estimate facial parameters from each frame of the driving video directly.
-Subsequently, we render these parameters into a sequence of 2D facial surface normal snapshots.
-However, in addition to containing the pose and expression information we require, these snapshots also include identity information reflected by the facial geometry from the driving frames.
-The identity information essential for our purpose should be extracted from the source image.
-Any inconsistency in the facial geometry information between the condition and the final output can result in sub-optimal performance, particularly noticeable when there is a significant mismatch in identity between the source image and the driving video.
-To address this issue, we utilize a facial shape alignment strategy which combines the identity information from the source image with the pose and expression information from the driving frames to generate the refined 2D facial conditions.
-Specifically, we obtain shape parameters $\beta$ containing identity information from the source image, and other parameters including pose $\theta$ and expression $\psi$ from the driving frames by the DECA encoder.
-We then use FLAME to generate the 3D face meshes and render them into surface normal snapshots to eliminate the effects of lighting conditions.
+Improved Spatial Conditioning with Facial Alignment
+We utilize the encoder of DECA to estimate head parameters from each frame of the target video directly.
+Subsequently, we render these parameters into a sequence of 2D surface normal snapshots for spatial-aligned conditioning.
+However, in addition to containing the pose and expression information we require, such snapshots inherently contain the identity characteristic reflected by the head geometry of the motion frame rather than that of the reference image.
+The inconsistency in the facial geometry between the spatial-aligned condition and the final output can result in sub-optimal performance, especially when there is a significant mismatch in identity characteristics between the reference image and the target video—a challenge that is particularly pronounced in stylized animation.
+Additionally, when there is a substantial head scale (the relative size or proportions of the face within the image) discrepancy between the reference image and the motion frames, the spatial information provided by the snapshots may conflict with that provided by the ReferenceNet.
+To address these issues, we utilize a strategy named Facial Alignment, which consists of two components: shape alignment, which integrates head shape from the reference image with pose and expression information from the target frames, and scale alignment, which ensures alignment of the 2D surface normal snapshots with the scale of the reference head.
+Specifically, we obtain shape parameters $\beta$ containing identity characteristic from the reference image, and other parameters including pose $\theta$ and expression $\psi$ from the target frames by the DECA encoder.
+The decoding process consists of generating 3D head vertices based on FLAME, applying the similarity transformation to head vertices using the rescaled transformation matrices, and render them into surface normal snapshots to eliminate the effects of lighting conditions.
 The entire process can be represented as:
-$n^{1:N}=\mathcal{R}(M(\beta_{s},\theta^{1:N}_{d},\psi^{1:N}_{d}),c^{1:N}_{d}),$ (2)
-where $n^{1:N}$ represents the 2D facial conditions, the rendered surface normal snapshots, for frames $1:N$.
+$n^{1:N}=\mathcal{R}(\mathcal{T}(\mathcal{F}(\beta_{s},\theta^{1:N}_{d},\psi^{1:N}_{d}),c^{1:N}_{d}),a^{1:N}),$ (2)
+where $n^{1:N}$ represents the 2D head conditions, the rendered surface normal snapshots, for frames $1:N$.
 $\mathcal{R}$ represents the rasterizer (PyTorch3D [58]).
-M represents FLAME model used to obtain head vertices with facial parameters.
+$\mathcal{F}$ represents FLAME model used to obtain head vertices with facial parameters.
 $c$ represents the camera information to project the 3D mesh into image space.
-$\beta_{s}$ is extracted from the source image, while $\theta^{1:N}_{d}$, $\psi^{1:N}_{d}$ and $c^{1:N}_{d}$ are extracted from the driving frames.
-After extracting the 2D facial conditions which are spatial-aligned with the output frames, we encode them with the Pose Encoder, a series of convolution layers, similar to Animate Anyone [53], to project the conditions to the same resolution with the noisy input of denoising UNet.
-Subsequently, we concatenate the spatial-aligned condition with the noisy input and feed it to the denoising UNet.
+$\beta_{s}$ is extracted from the reference input, while $\theta^{1:N}_{d}$, $\psi^{1:N}_{d}$ and $c^{1:N}_{d}$ are extracted from the target motion frames.
+$\mathcal{T}$ represents the similarity transformation with the rescaled matrices $a^{1:N}$.
+We derive a rescaled similarity transformation matrix for each frame, using a rescaling factor computed from the facial bounding boxes of the reference image and the first motion frame
+After extracting the 2D surface normal snapshots which are spatial-aligned with the output frames, we encode them with the Pose Encoder, a series of convolution layers, similar to Animate Anyone [50], to project the conditions to the same resolution with the noisy input of denoising UNet.
+Subsequently, we add the spatial-aligned condition to the noisy input and feed it to the denoising UNet.
 Expression Adapter
-Crucially, the low resolution of the 3D face mesh leads to the loss of mid-frequency details related to expression.
+Crucially, the low resolution of 3D face meshes results in the loss of mid-frequency expression details, which are essential for producing high-quality expressive stylized animations.
 Drawing inspiration from IP-Adapter [43] for image prompt conditioning, we introduce an expression adapter conditioning mechanism to address potential information loss resulting from spatial modulation, as illustrated in the lower branch of Fig.
-1.
-Unlike IP-Adapter for conditional image generation, we integrate video-level expression prompts from the driving frames in parallel with the image prompt from the source image.
-Specifically, we extract the expression embedding $\mathbf{\psi}\in\mathbb{R}^{N\times 50}$ from N driving frames using DECA and project it into the same dimension as the CLIP image embedding.
-This projected expression embedding $\mathit{p_{\psi}}$ is then injected into the diffusion model through additional cross-attention layers, which can be represented as:
-$Z_{new}=Attention(Q,K^{i},V^{i})+\lambda\cdot Attention(Q,K^{\psi},V^{\psi}),$ (3)
-where $\mathit{Q}$, $\mathit{K^{i}}$, and $\mathit{V^{i}}$ are the query, key, and value matrices from the original cross-attention layer for CLIP image features, while $\mathit{K^{\psi}}$ and $\mathit{V^{\psi}}$ are from the additional cross-attention layer for expression features.
-The projected expression embedding $\mathit{p_{\psi}}$ is injected into the cross-attention layer as $\mathit{K^{\psi}=p_{\psi}W_{k}^{\psi}}$ and $\mathit{V^{\psi}=p_{\psi}W_{v}^{\psi}}$, where $\mathit{W_{k}^{\psi}}$ and $\mathit{W_{v}^{\psi}}$ are the corresponding weight matrices.
-In the 3D denoising UNet for generating a video, the CLIP embedding of the source image is repeated N times to match the length of the expression embedding from the driving frames.
+2.
+Unlike IP-Adapter, which is used for conditional image generation, we incorporate video-level expression prompts (DECA expression embeddings) from the target motion frames in parallel with the image prompt (CLIP embedding) from the reference image.
+This module is designed to enhance the transfer of expressions while preserving the semantic information from the reference image.
+Specifically, we extract the expression embedding $\mathbf{\psi}^{1:N}\in\mathbb{R}^{50}$ from N motion frames using DECA and project it into the same dimension as the CLIP reference embedding.
+This projected expression embedding $\mathit{p_{\psi}^{1:N}}$ is then injected into the diffusion model through additional cross-attention layers, which can be represented as:
+$Z_{new}=Attention(Q,K_{i},V_{i})+\lambda\cdot Attention(Q,K_{\psi},V_{\psi}),$ (3)
+where $\mathit{Q}$, $\mathit{K_{i}}$, and $\mathit{V_{i}}$ are the query, key, and value matrices from the cross-attention layers for CLIP image features, while $\mathit{K_{\psi}}$ and $\mathit{V_{\psi}}$ are from the additional cross-attention layer for expression features.
+The projected expression embedding $\mathit{p_{\psi}^{1:N}}$ is injected into the cross-attention layer as $\mathit{K_{\psi}=p_{\psi}^{1:N}W^{k}_{\psi}}$ and $\mathit{V_{\psi}=p_{\psi}^{1:N}W^{v}_{\psi}}$, where $\mathit{W_{k}^{\psi}}$ and $\mathit{W_{v}^{\psi}}$ are the corresponding weight matrices.
+In the 3D denoising UNet for generating a video, the CLIP embedding of the reference image is repeated N times to match the length of the expression embedding from the motion frames.
 3.4 Training and Inference
 The training process comprises two stages.
-In the first stage, the model is trained to generate a single image under the guidance of a source image and a driving frame.
-During the training process, the model utilizes same-identity construction as its objective, where both the source image and the driving frame are derived from the same video.
-DECA is utilized to extract facial parameters and render facial snapshots from the driving frame.
-The driving frame $\textbf{x}_{0}$ is first encoded into the latent representation $\textbf{z}_{0}$, which is then noised with $\epsilon\sim\mathcal{N}(0,1)$ at timestep t via a defined scheduler as:
+In the first stage, the model is trained to generate a single image under the guidance of a reference image and a target motion frame.
+During the training process, the model utilizes same-identity construction as its objective, where both the reference image and the motion frame are derived from the same video.
+DECA is utilized to extract head parameters and render snapshots from the motion frame.
+The target motion frame $\textbf{x}_{0}$ is first encoded into the latent representation $\textbf{z}_{0}$, which is then noised with $\epsilon\sim\mathcal{N}(0,1)$ at timestep t via a defined scheduler as:
 $\textbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\textbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ (4)
 where $\bar{\alpha}_{t}={\textstyle\prod_{s=0}^{t}}\alpha_{s}$, $\alpha_{t}$ is the variance schedule, and $\textbf{z}_{t}$ is the noisy latent.
-The model is trained to predict the added noise by denoising function $\epsilon_{\theta}$ with condition features f from the source image and the driving video.
+The model is trained to predict the added noise by denoising function $\epsilon_{\theta}$ with condition features f from the reference image and the target video.
 The VAE, CLIP image encoder, and DECA are pretrained and kept frozen while the pose encoder, the expression adapter, the ReferenceNet, and the denoising UNet are updated.
@@
 $\textbf{L}_{stage1}=\mathbb{E}_{\textit{t},\textbf{f},\epsilon,\textbf{z}_{t}}[\|\epsilon-\epsilon_{\theta}(\textbf{z}_{t},\textit{t},\textbf{f})\|^{2}_{2}]$ (5)
-In the second stage, the models are trained to generate a temporally consistent video clip (N frames) with guidance from both a source image and a driving video clip.
+In the second stage, the models are trained to generate a temporally consistent video clip (N frames) with guidance from both a reference image and a target motion video clip.
 Pretrained temporal modules [57] are integrated into the denoising UNet.
@@
 $\textbf{L}_{stage2}=\mathbb{E}_{\textit{t},\textbf{f}^{1:N},\epsilon,\textbf{z}_{t}^{1:N}}[\|\epsilon-\epsilon_{\theta}(\textbf{z}_{t}^{1:N},\textit{t},\textbf{f}^{1:N})\|^{2}_{2}]$ (6)
-At inference, we utilize DECA to extract information from both the source image and the driving video.
-Leveraging a facial shape alignment strategy, we refine the facial snapshot sequence.
-Begining with noises $\textbf{z}_{t}^{1:N}$ sampled from a Gaussian distribution in the latent space, the model conduct denoising process to $\textbf{z}_{t-1}^{1:N}$ by predicting noise using the denoising UNet with the condition $\textbf{f}^{1:N}$ and iteratively to $\textbf{z}_{0}^{1:N}$ via a defined sampling process.
+At inference, we utilize DECA to extract information from both the reference image and the target video.
+Leveraging the facial alignment strategy, we refine the surface normal snapshot sequence.
+Beginning with noises $\textbf{z}_{t}^{1:N}$ sampled from a Gaussian distribution in the latent space, the model conduct denoising process to $\textbf{z}_{t-1}^{1:N}$ by predicting noise using the denoising UNet with the condition $\textbf{f}^{1:N}$ and iteratively to $\textbf{z}_{0}^{1:N}$ via a defined sampling process.
 Finally, the denoised outputs $\textbf{z}_{0}^{1:N}$ are projected back to the pixel level $\textbf{x}_{0}^{1:N}$ to obtain the generated video via the VAE decoder.
 4 Experiments
-In our experiments, we utilized the VoxCeleb dataset [2] after preprocessing as described in [5].
-This dataset comprises 19,522 and 525 utterance sequences, spanning 453 and 13 identities, with 3,354 and 90 videos in the training and test sets, respectively.
-Training was conducted on $256\times 256$ images using 2 NVIDIA A100 GPUs.
+4.1 Data and Training Details
+In our experiments, we generated a diverse array of styles (e.g.
+artistic, cartoon, sculpture, and painting) using [2] to conduct out-of-domain animation experiments.
+This broad spectrum of styles enabled us to evaluate the robustness and adaptability of our method across a variety of domains beyond the original dataset distribution.
+We utilized the subset of VoxCeleb dataset [59] after preprocessing as described in [5] for training and test.
+We also test the methods on the test set of VoxCeleb2 dataset [60].
+Training was conducted on $256\times 256$ images using 2 NVIDIA A100 80G GPUs for around 4 days.
 In the first stage, we trained for over 100,000 steps with a batch size of 32.
@@
 All models were optimized using the Adam optimizer with a learning rate of $1\times 10^{-5}$.
-4.1 Comparisons
+During the inference phase, our method generates results based on a given reference image and a target motion video, utilizing the DDIM sampler [61] with 20 denoising steps.
+The entire process takes approximately 25 seconds to produce a 100-frame video clip using a single NVIDIA A100 80G GPU.
+4.2 Out-of-Domain Animation
+Figure 3: Qualitative comparison with SOTA methods on out-of-domain stylized animation. Our method can generate high-quality images even without explicit training on such types of data.
+Figure 4: Qualitative comparison with SOTA methods on out-of-domain stylized animation in the presence of significant pose variations between reference images and target motion.
+Our method accurately captures the reference features and faithfully realize the target pose and expression, even under significant pose variations.
 We compare our method with four state-of-the-art methods, namely FOMM [5], Face vid2vid [6], TPSMM [7], and HyperReenact [26].
 Results for the baseline methods are generated using their provided code and pretrained weights, with the exception of Face vid2vid, for which we utilize unofficial code [^1] https://github.com/zhanglonghao1992/One-Shot_Free-View_Neural_Talking_Head_Synthesis due to the absence of the official version.
-We assess our method and the baseline approaches using four distinct metrics.
-First, we utilize the Fréchet-Inception Distance (FID) to evaluate image quality by measuring the distribution discrepancy between the generated data and the test set of VoxCeleb.
-Second, we employ the identity preservation cosine similarity (CSIM), following the approach outlined in Encoding in Style [61], to quantify the preservation of identity.
-Third, we measure the accuracy of pose and expression transfer using the Average Pose Distance (APD) and the Average Expression Distance (AED).
-CSIM is computed based on the source image and the outputs, while APD and AED are calculated based on the driving video and the outputs.
-Quantitative Comparisons
+We compare our method with state-of-the-art (SOTA) methods on out-of-domain data to assess its generalization capabilities across a diverse range of styles.
+As shown in Fig.
+3, our method demonstrates strong generalization on out-of-domain data, successfully animating characters from various styles and domains.
+Notably, stylized avatars can be animated using human-like virtual faces (as shown in row 1 and row 2 of both Fig.
+3 and Fig.
+4) as well as stylized avatars from other domains.
+Our method preserves the appearance and intricate details of the reference image across different styles while accurately capturing and generating the pose and expression from the motion video even without being trained on such data.
+In contrast, warping-based methods (FOMM, Face vid2vid, and TPSMM) exhibit noticeable identity leakage when there is a significant head shape difference between the reference and target images.
+HyperReenact, on the other hand, fails to preserve both the style and appearance of the reference image.
+This is likely due to the limitations of GAN inversion and the use of a pre-trained ArcFace model [62].
+This strong generalization capability makes our method highly adaptable for various applications, including animation production, privacy protection, virtual education, and beyond.
+It enables the animation of virtual avatars in diverse styles, offering flexibility and enhanced user experience.
+Furthermore, the ability to handle out-of-domain data addresses privacy and ethical concerns, as stylized or non-photorealistic avatars reduce the potential risks of misinterpretation or misuse.
+Fig.
+4 illustrates a more challenging scenario involving significant pose differences between the reference and target images.
+In this case, FOMM, Face vid2vid, and TPSMM tend to produce blurry outputs with severe artifacts and distortions due to their reliance on optical flow fields.
+The blurriness is particularly noticeable in regions where motion changes occur.
+Moreover, these methods struggle with inpainting, resulting in blurry and unnatural regions when generating areas not present in the reference image.
+HyperReenact, which does not depend on optical flow fields, generates finer-grained results compared to the warping-based methods.
+However, it still suffers from a loss of style, appearance, and fine details, such as hair texture and lighting.
+In contrast, our method excels in preserving detailed information from the reference input while accurately maintaining pose and expression from the target motion video, thanks to the specially designed conditioning module.
+Expression-related details, such as wrinkles, brow furrows, and lip curvature, can be faithfully generated.
+Overall, our method produces the cleanest, most high-fidelity results among all the state-of-the-art methods, even when only the side profile is available (as shown in row 1) or the head is partially occluded (as shown in row 4).
+4.3 Within-Domain Animation
+We assess our method and the baseline approaches using five distinct metrics on within-domain animation.
+First, we utilize the Fréchet-Inception Distance (FID) to evaluate image quality by measuring the distribution discrepancy between the generated data and the test set and the Fréchet Video Distance (FVD) to evaluate video quality.
+Both FID and FVD implementations are based on StyleGAN-V [63].
+Second, we employ the identity preservation cosine similarity (CSIM), following the approach outlined in Encoding in Style [64].
+Third, we measure the accuracy of pose and expression using the Average Pose Distance (APD) and the Average Expression Distance (AED).
+CSIM is computed based on the reference image and the outputs, while APD and AED are calculated based on the target video and the outputs.
 Table 1: Quantitative results on VoxCeleb. The best results are shown in bold, and the second-best results are shown in underline
 .
-Same-identity Reconstruction Cross-identity Reenactment FID CSIM APD AED FID CSIM APD AED User Pref.
-(%) FOMM [5] 21.91 0.83 0.011 0.092 41.93 0.51 0.028 0.210 10.25 Face vide2vid [6] 18.57 0.84 0.012 0.092 35.21 0.56 0.034 0.238 9.50 TPSMM [7] 20.18 0.84 0.009 0.081 38.91 0.52 0.024 0.191 16.50 HyperReenact [15] 72.72 0.48 0.011 0.098 73.35 0.40 0.017 0.181 20.75 Ours 17.21 0.75 0.011 0.091 32.58 0.57 0.022 0.174 43.00
+Same-identity Reconstruction Cross-identity Animation FID CSIM APD AED FVD FID CSIM APD AED FVD FOMM [5] 26.48 0.83 0.011 0.092 142.18 50.36 0.51 0.028 0.210 254.60 Face vide2vid [6] 23.43 0.84 0.012 0.092 136.46 43.23 0.56 0.034 0.238 210.34 TPSMM [7] 25.44 0.84 0.009 0.081 146.98 44.66 0.52 0.024 0.191 253.44 HyperReenact [26] 98.33 0.48 0.011 0.098 379.07 98.92 0.40 0.017 0.181 364.83 Ours 21.25 0.75 0.011 0.091 168.76 36.26 0.57 0.022 0.174 241.83
+Table 2: Quantitative results on VoxCeleb2. The best results are shown in bold, and the second-best results are shown in underline
+.
+Same-identity Reconstruction Cross-identity Animation FID CSIM APD AED FVD FID CSIM APD AED FVD FOMM [5] 43.53 0.77 0.019 0.124 178.42 59.67 0.47 0.035 0.235 242.93 Face vide2vid [6] 40.91 0.78 0.021 0.124 181.64 51.13 0.54 0.039 0.232 238.19 TPSMM [7] 45.52 0.76 0.018 0.109 196.89 53.39 0.48 0.032 0.231 251.45 HyperReenact [26] 101.07 0.48 0.013 0.101 380.31 97.72 0.42 0.016 0.163 405.88 Ours 34.46 0.67 0.016 0.119 185.54 33.81 0.55 0.024 0.188 233.91
 We evaluate our method and the baselines on the test set of VoxCeleb.
 In the evaluation, higher scores indicate better performance for CSIM, while lower scores are preferable for all other metrics.
-We conduct one-shot face reenactment on two tasks: same-identity reconstruction, which mirrors the training process, and the more challenging primary task of cross-identity reenactment.
-For the same-identity reconstruction task, we randomly selected 100 videos from the VoxCeleb test set, using the first frame as the source image and considering the remaining frames as the driving frames.
+We conduct one-shot animation on two tasks: same-identity reconstruction, which mirrors the training process, and the more challenging primary task of cross-identity animation.
+For the same-identity reconstruction task, we randomly selected 100 videos from the VoxCeleb test set, using the first frame as the reference image and considering the remaining frames as the motion frames.
 As shown in Table 1, our method achieves the best image quality (FID) and ranks second for pose (APD) and expression accuracy (AED).
 TPSMM outperforms all other methods in terms of CSIM, APD, and AED.
-All warping-based methods (FOMM, Face vid2vid, and TPSMM) perform well on this proxy task when there is no identity variance and usually a small pose difference between the source image and the driving video.
-Even though current methods perform well on same-identity reconstruction, the main task (cross-identity reenactment), where the source image and driving video are from different identities, introduces substantial identity mismatches and pose variations, demanding higher generalization capabilities from models.
+All warping-based methods (FOMM, Face vid2vid, and TPSMM) perform well on this proxy task when there is no appearance variance and usually a small pose difference between the reference image and the target motion.
+Even though current methods perform well on same-identity reconstruction, the main task (cross-identity animation) introduces substantial appearance mismatches and pose variations, demanding higher generalization capabilities from models.
 Similar to previous work, we randomly select 100 image-video pairs from the VoxCeleb test set.
-Table.1 illustrates a significant performance drop for each method when transitioning from same-identity reconstruction to cross-identity reenactment.
+Table.1 illustrates a significant performance drop for each method when transitioning from same-identity reconstruction to cross-identity animation.
 As observed in previous SOTA methods, there is a trade-off between identity consistency (CSIM) and expression accuracy (AED).
-The more modifications related to expression made to the original source identity, the less the generated image will resemble the original, as assessed by face recognition models.
+The more modifications related to expression made to the original reference image, the less the generated identity characteristics will resemble the original, as assessed by facial models.
 Despite this trade-off, Our method (AniFaceDiff) achieves SOTA results on both CSIM and AED.
-Additionally, we outperform all four baseline methods in terms of image quality (FID) and secure the second position in terms of pose accuracy (APD).
-Qualitative Comparisons
-Figure 2: Qualitative comparison with SOTA methods on cross-identity reenactment.
-Our method achieves high identity preservation and expression accuracy.
-(All photorealistic human images presented in this paper are virtual individuals generated by [33] and do not exist in the real world.)
-Figure 3: Qualitative comparison with SOTA methods on cross-identity reenactment in the presence of significant pose variations between the source and driving images.
-Our method accurately captures the source identity and faithfully transfers the target pose and expression, even under significant pose variations.
-We present a few representative examples of cross-identity reenactment in Fig.
-2.
-Additional examples can be found in Appendix A.2.
-Overall, our method generates high quality images compared to all other state-of-the-art methods while preserving both source identity and pose and expression of the driving frame.
-Better preservation of source identity is especially evident in the first and second rows, where other methods show noticeable influence from the driving identity.
-In contrast, our method maintains the source identity with greater accuracy.
-Furthermore, our approach excels at maintaining the expressions and poses of the driving videos, as seen in the third and fourth rows, where our method consistently captures expression-related details such as wrinkles, brow furrows and lip curvature more accurately than other state-of-the-art methods.
-Fig.3 demonstrates a more challenging scenario where there are significant pose differences between the source and driving images.
-FOMM, Face vid2vid, and TPSMM tend to produce blurry outputs with severe artifacts and distortions due to their reliance on the optical flow field.
-Additionally, their inadequate inpainting capabilities result in blurry and unrealistic regions when generating areas not present in the source image.
-HyperReenact, which does not rely on the optical flow field, generates more fine-grained results.
-However, it still suffers from a loss of identity and detailed information, such as facial wrinkles, hair texture, and lighting.
-This is likely due to the limitations of GAN inversion and the use of a pretrained ArcFace [62] for extracting identity information from the source image.
-In contrast, our method excels in preserving detailed information from the source image, while accurately maintaining pose and expression from the driving video due to the designed conditioning module.
-Overall, our method produces the cleanest and most high-fidelity results among all the state-of-the-art methods, even when the source face is partially occluded (row 3) or when there is a large distance difference from the camera between the source and driving images (row 4).
-We note that the current metrics might not fully capture the qualitative differences observed in the results.
-Hence, we conducted a user study involving 20 participants to further assess performance based on human visual experience.
-We presented 20 randomly selected source-driving image pairs: 5 for same-identity reconstruction and 15 for cross-identity reenactment, as cross-identity reenactment is the primary task.
-We provided users with three evaluation criteria similar to those used by HyperReenact [15]: 1) image quality, 2) identity and nuanced detail preservation from the source image, and 3) pose and expression preservation from the driving image.
-As shown in Table 1, our method significantly outperforms all other SOTA methods based on user preference.
-Out-of-domain Data
-Figure 4: Qualitative comparison with SOTA methods on out-of-domain (non-photorealistic) face reenactment.
-Our method can generate high-quality images even without being trained on such types of data.
-(All out-of-domain facial images are generated by [63])
-We also compare our method with these SOTA methods outside the training distribution to evaluate generalization capabilities.
-As shown in Fig.
-4, The first and second rows demonstrate out-of-domain (non-photorealistic) source images reenacted by out-of-domain driving images.
-The third and fourth rows demonstrate out-of-domain source images reenacted by in-domain (photorealistic) driving images.
-Our method exhibits strong generalization capabilities on out-of-domain data.
-More results can be found in Appendix A.3.
+Additionally, our method outperforms all four baseline methods in terms of image quality (FID) and secures the second position in terms of pose accuracy (APD) and video quality (FVD).
+To assess the generalization ability, we tested both our method and other state-of-the-art approaches on the VoxCeleb2 test set (all models were trained on VoxCeleb).
+Similar to the evaluation protocol used for VoxCeleb, for same-identity reconstruction, we randomly selected 100 videos from the VoxCeleb2 test set, using the first frame as the reference image and the remaining frames as motion frames.
+For cross-identity animation, we randomly selected 100 pairs from the test set.
+The results are presented in Table 2.
+Consistent with the VoxCeleb evaluation, our method dominates in FID for same-identity reconstruction.
+For cross-identity animation, it significantly outperforms other approaches in FID and achieves the top scores in both CSIM and FVD, demonstrating strong generalization ability.
+Additionally, our method ranks second in APD and AED.
 4.4 Ablation Study
-Table 2: Quantitative comparison with or without Facial Shape Alignment and the Expression Adapter.
-The addition of both components allows for a better trade-off between identity consistency vs expression fidelity.
+Table 3: Quantitative comparison with or without Facial Alignment and the Expression Adapter.
+The addition of both components allows for a better trade-off between identity characteristics consistency vs expression fidelity.
 We find that combining both provides the best qualitative and quantitative results.
 .
-Same-identity Reconstruction Cross-identity Reenactment FSA EA CSIM APD AED CSIM APD AED - - 0.75 0.011 0.098 0.52 0.021 0.182 - - - - 0.59 0.022 0.206 - 0.75 0.011 0.091 0.51 0.021 0.158 - - - 0.57 0.022 0.174
-Figure 5: Qualitative comparison on cross-identity reenactment with or without Facial Shape Alignment and the Expression Adapter.
-FSA improves facial generation quality, while EA maintains expression consistency.
+Same-identity Reconstruction Cross-identity Animation FA EA CSIM APD AED CSIM APD AED - - 0.75 0.011 0.098 0.52 0.021 0.182 - - - - 0.59 0.022 0.206 - 0.75 0.011 0.091 0.51 0.021 0.158 - - - 0.57 0.022 0.174
+Figure 5: Qualitative comparison of results with or without Facial Alignment and the Expression Adapter.
+FA improves the generation quality, while EA maintains expression consistency.
 Ensembling both provides the best overall consistency.
 These observations are consistent with the quantitative results in Table 2.
-We conduct an ablation study on the proposed 1) Facial Shape Alignment (FSA) and 2) Expression Adapter (EA).
-Facial Shape Alignment is used exclusively for cross-identity reenactment.
-Our baseline method directly conditions the model with 2D facial snapshots extracted from the driving video, without the facial shape alignment and the expression adapter.
-As illustrated in Table 2, adding facial shape alignment significantly improves identity preservation (CSIM).
-Facial shape alignment uses the shape embedding from the source image to render the 2D facial snapshots, thereby enhancing identity preservation.
-Without facial shape alignment, the output maintains the shape from the driving image rather than the source image, as shown in Fig.
+We conduct an ablation study on the proposed 1) Facial Alignment (FA) and 2) Expression Adapter (EA).
+Facial Alignment is used exclusively for cross-identity animation.
+The baseline method, without FA and EA, is similar to the structure of AnimateAnyone, incorporating 2D surface normal snapshots from DECA.
+As illustrated in Table 3, adding FA significantly improves identity characteristics preservation (CSIM).
+FA uses the shape embedding from the reference image to render the 2D surface normal snapshots, thereby enhancing appearance preservation.
+Without it, the output tends to adopt the shape of the target motion rather than preserving the reference image’s features.
+This is particularly problematic when the reference and target images are from different styles, as it exacerbates the appearance mismatch, as shown in Fig.
 5.
-The expression adapter significantly enhances expression accuracy (AED) in both same-identity reconstruction and cross-identity reenactment.
-Methods incorporating the expression adapter can provide more expression-related details, such as wrinkles, and achieve greater accuracy in certain areas, such as the mouth, as depicted in Fig.
+On the other hand, FA ensures that the spatial scale of the surface normal snapshots matches that of the ReferenceNet, enhancing the spatial consistency of the generated output, especially when there is a substantial facial scale difference between the reference image and target motion.
+EA significantly enhances expression accuracy (AED) in both same-identity reconstruction and cross-identity animation.
+Methods incorporating EA can provide more expression-related details, and achieve greater accuracy in certain areas, such as the mouth and eyes, as depicted in Fig.
 5.
-Combining both FSA and EA can effectively handle the trade-off between identity consistency and expression accuracy.
+Combining both FA and EA can effectively handle the trade-off between appearance consistency and expression accuracy.
 5 Ethical Considerations
-We are committed to positive virtual face reenactment applications.
-All photorealistic human portraits presented in this paper are virtual and do not exist in reality.
-Acknowledging the broad applicability of face reenactment, we are also aware of its potential for malicious misuse, such as in deepfake fraud.
-To address this concern, we plan to implement stringent access control mechanisms to limit access permissions exclusively for research purposes.
-Additionally, we evaluate existing SOTA deepfake detectors (see Appendix A.1) and conduct an analysis to identify potential shortcomings.
+We are committed to the positive applications of virtual stylized avatar animation in fields such as film, video conferencing, virtual education, and telemedicine.
+For instance, it can be used to safeguard privacy [65], protecting real identities in various digital environments, including social media, live streaming, telemedicine, and sensitive interviews.
+Acknowledging the broad applicability of these technologies, we are also aware of their potential for misuse when combined with other methods on photorealistic heads, as they cannot independently create complete videos with audio.
+We firmly oppose the creation of misleading or harmful content and aim to help promote the ethical use of these technologies.
+We recommend that all applications of these technologies incorporate watermarks into the images they produce.
+Furthermore, detection for generative content provides a practical solution to this issue, serving as an effective defender by identifying whether a video is real or generated.
+We are dedicated to contributing to this field by supporting the development of detectors.
+To this end, we evaluate the effectiveness of current state-of-the-art detectors and analyze their generalization capabilities, helping to identify potential areas for improvement and suggesting possible solutions in 5.1.
 This will serve as a foundation for future research aimed at mitigating weaknesses of existing detectors and providing benchmarking data for future approaches.
-Appendix A Appendix / supplemental material
-A.1 On deepfake detection and potential for misuse
-While any face reenactment technique can also be misused for deepfake generation, it is clearly an important area of active research with many ethical applications in movies, journalism, animation and virtual reality, to name a few.
-This is especially apparent in emergent capabilities in techniques such as ours, as improved results in the primary domain (real human face animation) can lead to significant improvements in cross-domain or out of domain animation transfer (human to 3D, human to cartoon, cartoon to 3D, cartoon to cartoon etc.) However, to mitigate the deepfake generation potential of our approach, we perform a detailed and comprehensive analysis of state of the art detection techniques on our method, and report their zero shot performance on our data.
-First, this contributes to the understanding of the generalization capabilities of existing deepfake detectors.
-Secondly, it highlights potential for improving the training datasets of such detectors, especially with videos based on paradigms using diffusion models as the generative core.
-This work opens up the possibility of creating a large scale dataset of high fidelity diffusion based face reenactment videos, addressing the issue of lack of such data in the deepfake detection domain.
-Deepfake Autoencoder (DFAE, e.g.
-DeepFaceLab [^2] https://github.com/iperov/DeepFaceLab) and other face swapping methods can produce highly realistic videos that are challenging to detect using existing detection algorithms.
-This is evident in the results of detection algorithms on the DFDC [64] dataset.
-Additionally, videos generated by face reenactment methods such as Neural Texture [65] and Face2Face [66] are also difficult to detect using off-the-shelf detectors (details in [67]).
-It is worth noting that these methods do not offer any alternative means of detecting videos generated by them, which encourages generation of defamatory and harmful deepfakes.
+Additionally, we are committed to strictly limiting our model’s use to ethical research purposes and actively supporting the detection community by sharing insights and resources.
+Privacy considerations can be found in 5.2.
+As responsible researchers, we call on other scholars to prioritize ethical considerations in their work.
+Additionally, our current method cannot produce videos with photorealistic avatars that resemble real-life footage without visible artifacts and flickering issues.
+5.1 Generalization Analysis in Synthetic Face Detection
+Stylized avatar animation is an important area of active research with numerous ethical applications.
+However, as with any generative technology, there are inherent risks related to potential misuse.
+Detection for synthetic face serves as a important safeguard by identifying generated content and distinguishing it from authentic footage, thus helping prevent the spread of misinformation and protecting individuals from potential harm.
+To address this risk, we conduct a detailed and comprehensive analysis of state-of-the-art detection techniques on our method and report their zero-shot performance results.
+This contributes to the understanding of the generalization capabilities of existing detectors, which is essential for advancing responsible applications.
+Table 4: Detection frame-level and video-level AUC from the state-of-the-art detectors with respective pretrained weights on videos generated by our method AniFaceDiff.
+The best results are shown in bold.
 Model Frame-AUC Video-AUC
@@
 Average 84.92 86.52
-Table 3: Detection frame-level and video-level AUC from the state-of-the-art deepfake detectors with respective pretrained weights on videos generated by our reenactment method AniFaceDiff.
-We emphasize the significance of ensuring that these methods are used ethically and that they can be detected with existing deepfake detectors; this would invalidate videos generated with malicious intent.
-We evaluated videos generated with our method on pretrained, state-of-the-art deepfake detection methods.
+We emphasize the importance of using generative technologies responsibly, ensuring that generated content can be effectively detected.
+To this end, we evaluated the videos produced by our method using pretrained, state-of-the-art detection techniques.
 Table 4 presents both frame and video-level detection results.
-Although video-level detection is more common, we include frame-level detection due to the possibility of temporally partial deepfakes [67] i.e.
-having both real and fake segments in the same video.
+Although video-level detection is more common, we include frame-level detection due to the possibility of temporally partial synthesized videos [74] i.e.
+having both real and generated segments in the same video.
 Our evaluation encompasses multiple types of detectors including naive detectors [66, 68], a frequency-level detector [70] and spatial detectors [67, 69, 71, 72, 73].
@@
 Frame-level detection is more challenging since the detection performance depends on the prediction of each frame, rather than a majority voting across the frames in a video.
-While the results are promising, there is room for improvement.
-Further investigation into diffusion-based reenactment methods could further enhance detection rates.
+Among the recent detection methods tested, CORE and CADDM achieve higher AUC scores, likely due to their enhanced feature learning capabilities.
+These results align with the specific features each method targets.
+CORE employs explicit regularization to ensure representation consistency across different augmentations, effectively mitigating overfitting.
+On the other hand, CADDM addresses implicit identity leakage, thereby improving its generalization on unseen data.
+Overall, enhancing the generalization capabilities of detection algorithms is crucial to mitigating overfitting and ensuring their effectiveness across different generative methods.
+Different characteristics from generative methods sometimes make it challenging for a single detector to perform optimally across all scenarios.
+By utilizing multiple detectors that employ different feature-learning strategies, we can better address the challenges posed by synthetic content and improve the reliability of detection, thus providing more comprehensive and robust safeguards against potential misuse.
 This assessment aims to pave the way for future studies on detection.
-Additionally, we provide access to researchers from deepfake detection in order to facilitate the generation of higher quality deepfake datasets, particularly those generated using state of the art face reenactment diffusion models (such as our work) - a current gap in the deepfake literature.
-A.2 More results on cross-identity reenactment
-Figure 6: Additional qualitative comparison with SOTA methods on cross-identity reenactment.
-A.3 More results on out-of-domain reenactment
-Figure 7: Additional qualitative comparison with SOTA methods on out-of-domain reenactment.
-A.4 Limitations
-Although our method provides high-fidelity results with excellent image quality, there are still some limitations.
-First, noticeable flickering occurs when generating videos.
-Second, the inadvertent generation of hands affects the quality of the generated content.
-Future studies can focus on addressing these issues.
+Additionally, we aim to support the creation of higher-quality datasets for detection, particularly those generated using state-of-the-art diffusion models-based methods, addressing a current gap in the literature.
+5.2 Privacy Considerations
+Dataset Usage
+This research utilizes the VoxCeleb and VoxCeleb2 datasets developed by the Visual Geometry Group (VGG) at the Department of Engineering Science, Oxford University, which come with an explicit privacy notice [^2] https://www.robots.ox.ac.uk/~vgg/terms/url-lists-privacy-notice.html.
+We use these datasets to evaluate the model, conduct benchmarking comparisons with previous methods, and assess detectors.
+Notably, we do not use any metadata, such as identity or nationality, in this research.
+Protection of Privacy and Data Integrity
+In our research, we recognize the importance of ethical considerations and privacy concerns.
+Our focus is on animating stylized avatars, which can be utilized to protect privacy in online environments.
+To address privacy concerns, we take several measures: First, all characters presented in this paper are entirely virtual and do not correspond to real individuals.
+Second, this study does not aim to identify or analyze personal data within publicly available datasets in any way that could lead to re-identification.
+We minimize the personal data by removing personal identity information (metadata) and voice data from the dataset, further preventing cross-correlation between data sources.
+Third, all images in the datasets are resized to 256x256, effectively reducing facial details and enhancing de-identification.
+Finally, access to the data, models, and model-generated outputs are strictly controlled.
+No private or personally identifiable information from individuals is collected or used in this research.
+All data processing adheres to privacy standards and does not involve identity recognition.
+We advocate for the responsible use of these technologies, emphasizing that they are applied in ways that uphold individual rights and contribute positively to society.
+6 Conclusion
+In this work, we introduce Anifacediff, a novel framework built on the Stable Diffusion model specifically designed for stylized avatar animation.
+The objective is to accurately maintain reference image consistency while effectively animate it with the pose and expression derived from the target motion video.
+To achieve this, we developed a new conditioning module that integrates two key components: Facial Alignment (FA) and Expression Adapter (EA).
+Facial Alignment ensures that the identity characteristics of the reference image are preserved and that spatial information is consistently aligned, while the Expression Adapter captures and conveys the intricate emotional nuances present in the target motion.
+Extensive experiments demonstrate that our method outperforms state-of-the-art techniques, especially when handling out-of-domain data across a wide range of styles.
+Additionally, our analysis of current state-of-the-art detectors evaluates their generalization capabilities, highlighting potential areas for improvement and possible solutions.
+This contribution reinforces our commitment to advancing detection techniques for generative content and promoting responsible development.
+6.1 Limitations
+Although our method delivers high-quality results, it has several limitations.
+First, flickering occurs during video generation.
+Second, the inadvertent generation of hands negatively impacts the quality of the produced content.
+Furthermore, the model’s performance is compromised if the character to be animated is overly abstract.
+Future studies could focus on addressing these issues.
+Finally, the low inference speed of diffusion models hinders the application in real-time scenarios.
+This limitation could be mitigated by developing faster samplers for diffusion models while maintaining high-quality generation.
+Figure 6: Additional stylized animation results produced by our method.